OpenStudy (anonymous):

What is the magnitude and direction of line GH if G(2,-2) and H(7,6)?

OpenStudy (anonymous):

magnitude: 9.43 units; direction 32.01° magnitude: 6.4 units; direction: 57.99° magnitude: 9.43 units: direction: 57.99° magnitude: 6.4 units: direction: 32.01°

OpenStudy (lgbasallote):

\[\text{magnitude} = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\] \[\text{direction} = \tan^{-1} \left(\frac{y_2 - y_1}{x_2 - x_1} \right)\] does that help?

OpenStudy (anonymous):

i dont really understand this stuff at all:(

OpenStudy (anonymous):

i know i have to plug x and y to where it sais you do i just dont have the calculator with me lol